OpenStudy (anonymous):

. The box-and-whisker plot shown summarizes the number of laps team members swam in practice. What is the interquartile range for this data? A. 10 B. 20 C. 15 D. 5

OpenStudy (texaschic101):

Do you know what Q3 and Q1 are ? Because interquartile range (IQR) is Q3 - Q1

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i really need help

OpenStudy (texaschic101):

Q1...it is the beginning of the box...not the whiskers, just the box....so Q1 = 10 Q3 ..it is the end of the box..so it is 25 So the IQR is 25 - 10 = 15 understand ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

my answer is C

OpenStudy (texaschic101):

correct :)

OpenStudy (anonymous):

Which is a true statement concerning outliers for the data set summarized by the box-and-whisker plot shown? A. This data set has outliers at both extremes. B. The minimum value is 10 and it is an outlier. C. The maximum value is 80 and it is an outlier. D. The minimum value is 10, but it is not an outlier.

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

i really need help i am going to be late for a meeting so i really neeed help could you help me

OpenStudy (anonymous):

i think the answer is C

OpenStudy (anonymous):

am i correct

OpenStudy (texaschic101):

let me check... to find outliers...multiply IQR by 1.5 10 * 1.5 = 15 add to Q3....15 + 56 = 71 (anything above 71 is an outlier) subtract from Q1...46 - 15 = 31 (anything below 31 is an outlier) answer is A